### v1

# Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability

Abstract
We investigate Bayesian posterior consistency in the context of parametric models with proper priors. While typical state-of-the-art approaches rely on regularity conditions that are difficult to verify and often misaligned with the actual mechanisms driving posterior consistency, we propose an alternative framework centered on a simple yet general condition we call sequential identifiability. This concept strengthens the usual identifiability assumption by requiring that any sequence of parameters whose induced distributions converge to the true data-generating distribution must itself converge to the true parameter value. We demonstrate that sequential identifiability, combined with a standard Kullback–Leibler prior support condition, is sufficient to ensure posterior consistency. Moreover, we show that failure of this condition necessarily entails a specific and pathological form of oscillations of the model around the true density, which cannot exist without intentional design. This leads to the important insight that posterior inconsistency may be safely ruled out, except in the unrealistic scenario where the modeler possesses precise knowledge of the data-generating distribution and deliberately incorporates oscillatory pathologies into the model targeting the corresponding density. Taken together, these results provide a unified perspective on both consistency and inconsistency in parametric settings, significantly expanding the class of models for which posterior consistency can be rigorously established. To illustrate the strength and versatility of our framework, we construct a one-dimensional model that violates standard regularity conditions and fails to admit a consistent maximum likelihood estimator, yet supports a complete posterior consistency analysis via our approach.
1 Introduction
Asymptotic consistency is a fundamental criterion for evaluating the quality of statistical estimation procedures. In Bayesian inference, the study of consistency of posterior distributions has been a highly active area of research, with substantial recent progress particularly in nonparametric settings. A prominent example is the ubiquitous problem of density estimation from independently and identically distributed (iid) observations, where key contributions by Ghosal et al., (1999), Barron et al., (1999) and Walker, (2004) address consistency by sieve methods ensuring that the posterior distribution concentrates in any appropriately defined neighborhood of the true data-generating density.
In contrast, this article focuses on the study of posterior consistency in parametric models. To set the stage, consider the following standard modeling framework: let the sample space be $(\mathbb{R},\mathscr{B}(\mathbb{R}))$, where $\mathscr{B}(T)$ denotes the Borel $\sigma$-algebra on a topological space $T$, and let $\mathrm{d}x$ denote the Lebesgue measure on the real line. Define the statistical model $\mathcal{F}_{\Theta}:=\{f_{\theta}:\theta\in\Theta\}$, where $f_{\theta}:=\mathrm{d}F_{\theta}/\mathrm{d}x$ is the density (Radon–Nikodym derivative) associated with a probability measure [^1] With a slight abuse of notation, we identify any probability measure $F$ on $(\mathbb{R},\mathscr{B}(\mathbb{R}))$ with its cumulative disribution function (cdf), writing $F(x)\equiv F((-\infty,x])$ for all $x\in\mathbb{R}$. $F_{\theta}(\mathrm{d}x)\ll\mathrm{d}x$ parametrized by $\theta\in\Theta\subseteq\mathbb{R}^{p}$, for some $p\in\mathbb{N}$, and where $\Theta$ is assumed to be closed. As is natural for any sensibly parametrized family $\mathcal{F}_{\Theta}$ intended for density estimation, we assume that convergence of parameters in $\Theta$ is equivalent to convergence of the associated densities with respect to some strong metric, such as the Hellinger distance $d_{h}$; that is, for all $\theta,\theta_{1},\theta_{2},\ldots\in\Theta$,
$\lim_{k\to\infty}\|\theta_{k}-\theta\|=0\iff\lim_{k\to\infty}d_{h}(f_{\theta_{k}},f_{\theta})=0;$
see Section 2 for a formal definition of $d_{h}$.
[…]
$\Pi(\mathrm{d}\theta\mid X_{1:n})=\frac{\prod_{i=1}^{n}f_{\theta}(X_{i})\,\Pi(\mathrm{d}\theta)}{\int_{\Theta}\prod_{i=1}^{n}f_{\theta^{\prime}}(X_{i})\,\Pi(\mathrm{d}\theta^{\prime})}.$
Posterior consistency at $\theta_{\star}$ is then formulated as the requirement that
$\lim_{n\to\infty}\Pi(A_{\varepsilon}^{c}\mid X_{1:n})=0\quad\text{a.s.-}F_{\theta_{\star}}^{\infty}$ (1)
for all $\varepsilon>0$, where $A_{\varepsilon}:=\{\theta\in\Theta:\|\theta-\theta_{\star}\|<\varepsilon\}$. Given the equivalence between the Euclidean and Hellinger metrics on $\mathcal{F}_{\Theta}$, consistency implies that the posterior increasingly concentrates on densities that are arbitrarily close to the true density with respect to the Hellinger distance. Recalling that $d_{h}$ and the $L^{1}$ metric are themselves equivalent on $\mathcal{F}_{\Theta}$, this provides a meaningful notion of convergence for densities—namely, pointwise closeness integrated with respect to the Lebesgue measure.
Posterior consistency in parametric models was first studied in the foundational work of doob1949application, who showed that, under a mild asymptotic estimability condition, the posterior distribution is consistent at any $\theta_{\star}$ belonging to a set $\Theta_{\star}\subseteq\Theta$ of prior mass one. While remarkably parsimonious in its assumptions, Doob’s approach suffers from a major drawback: the set $\Theta_{\star}$ is not explicitly characterized by the theorem, making it impossible to verify whether a given $\theta_{\star}\in\Theta$ belongs to it. Moreover, the complement of $\Theta_{\star}$—where posterior inconsistency may occur—could, in principle, be large.
In contrast to Doob’s approach, the conditions currently considered state of the art for establishing parametric consistency rely on directly imposing smoothness and regularity assumptions on the model $\mathcal{F}_{\Theta}$. The work of Walker, (1969) (see also Berk,, 1970) is a prominent example that continues to be cited in contemporary applied and methodological literature on Bayesian methods (e.g., Mao et al.,, 2024; Rustand et al.,, 2023; Miller,, 2021; Doğan et al.,, 2021). Specifically, Walker, (1969) established posterior consistency by requiring the log-likelihood mapping $\theta\mapsto\ln f_{\theta}(x)$ to possess a sufficiently regular modulus of continuity and to exhibit suitable decay outside compact subsets of $\Theta$ (see Section 3 for a detailed discussion). Under these conditions, consistency of both the maximum likelihood estimator (MLE) and the posterior distribution follows. Although conditions of this type remain the gold standard for establishing parametric consistency (see, e.g., the recent textbook treatment in Ghosal and van der Vaart,, 2017, sect. 6.4), in this article we challenge their suitability by demonstrating that they are overly restrictive, misaligned with the fundamental mechanisms governing posterior convergence, and often difficult to verify in parametric models beyond the most basic cases.
1.1 Our contribution
To address this gap and develop a unified framework for the study of parametric posterior consistency, we advocate a foundational reexamination of the problem. Our starting point is the seminal result of Schwartz, (1964), who showed that if the model is identifiable and $\theta_{\star}$ lies in the Kullback–Leibler (KL) support of the prior $\Pi$ (see Section 2 for a precise definition), then the posterior is weakly consistent at $\theta_{\star}$, i.e., consistent with respect to the topology of weak convergence. This is equivalent to stating that the posterior concentrates on parameter values $\theta\in\Theta$ whose associated cdfs $F_{\theta}$ are close to $F_{\theta_{\star}}$, or, more precisely, those $F_{\theta}$ lying in any arbitrarily small weak neighborhood of $F_{\theta_{\star}}$.
The difficulty in translating weak consistency into parameter consistency lies in the fact that the standard assumption of identifiability does not account for weak limiting behavior. As a result, it remains possible for a sequence of cdfs in the model to converge to $F_{\theta_{\star}}$ while the corresponding sequence of parameters fails to converge to $\theta_{\star}$. This creates a gap between weak consistency and consistency in the sense of parameter values: under the assumptions of Schwartz’s theorem, the posterior may accumulate on regions of the parameter space that yield the correct distribution function but are not themselves close to the true parameter value. To address this loophole, our key contribution is to introduce a new notion of identifiability, which we term sequential identifiability, requiring that no sequence of parameters failing to converge to $\theta_{\star}$ can give rise to a sequence of cdfs converging to $F_{\theta_{\star}}$. In light of Schwartz’s theorem, and by incorporating weak limits into its formulation, sequential identifiability ensures that the posterior concentrates around the true parameter value and, consequently, the true density function.
We argue that, in parametric settings, sequential identifiability provides a fundamental yet non-invasive tightening of standard identifiability. Indeed, we show that it holds for many widely used parametric models and that violations leading to inconsistency can occur only when the model includes densities that oscillate arbitrarily often around the true density. Consequently, unless the modeler possesses precise knowledge of the data-generating process and deliberately incorporates specifically designed pathological oscillations—two highly unrealistic conditions in any practical parametric modeling scenario—the model can be declared sequentially identifiable and, by default, consistent. To illustrate these ideas, we construct a parametric model based on cosine oscillations that violates classical regularity conditions and fails to satisfy MLE consistency, yet exhibits posterior consistency at all but one parameter value, thanks to sequential identifiability. The extreme oscillatory behavior of this model, together with the fact that issues with sequential identifiability arise only at a single parameter value, illustrates our point: failures of posterior consistency can be almost universally ruled out in parametric models of practical relevance, unless one intentionally designs a model that violates sequential identifiability by knowing which true density to target with oscillations.
1.2 Summary of the article
The rest of the article is structured as follows. After establishing basic notation and definitions in Section 2, Section 3 reviews some classical conditions proposed in the literature to address posterior consistency in parametric models. In Section 4, we introduce our alternative approach based on sequential identifiability, providing general sufficient conditions for posterior consistency and linking inconsistency to a peculiar oscillatory behavior of the densities in the model. Section 5 illustrates our framework through a one-dimensional parametric model which, despite violating classical regularity conditions, is readily handled using our methodology. This example also serves to highlight key aspects of our theoretical analysis, particularly in relation to oscillations. Section 6 concludes the paper, while the proofs of all the theoretical results can be found in the supplementary material at the end of the article.
2 Notation and basic definitions
[…]
The Hellinger distance and the KL divergence between two densities $f$ and $g$ are defined respectively as
$d_{h}(f,g)$ $:=\left[\int_{\mathbb{R}}\left(\sqrt{f(x)}-\sqrt{g(x)}\right)^{2}\mathrm{d}x\right]^{1/2},$
$\textnormal{KL}(f,g)$ $:=\int_{\mathbb{R}}\ln\left(\frac{f(x)}{g(x)}\right)f(x)\,\mathrm{d}x,$
while the Lévy–Prokhorov metric between two probability measures $F$ and $G$ is given by
$d_{w}(F,G):=\inf\left\{\delta>0:\,\forall A\in\mathscr{B}(\mathbb{R}),\,F(A)\leq G(A^{\delta})+\delta,\,G(A)\leq F(A^{\delta})+\delta\right\},$
[…]
The next two definitions are fundamental. In particular, the KL support condition is central to Schwartz’s theorem (see Theorem 1 below) and forms the starting point of our analysis. As for identifiability (Casella and Berger,, 2024), we assume without further mention that all considered statistical models satisfy this basic property at all parameter values.
Definition 1.
A parameter $\theta_{\star}\in\Theta$ is said to be in the KL support of the prior $\Pi$, denoted $\theta_{\star}\in\textnormal{KLS}(\Pi)$, if
$\Pi\left(\left\{\theta\in\Theta:\textnormal{KL}(f_{\theta_{\star}},f_{\theta})<\varepsilon\right\}\right)>0\quad\text{for all }\varepsilon>0.$
Definition 2.
The parametric family $\mathcal{F}_{\Theta}:=\{f_{\theta}:\theta\in\Theta\}$ is identifiable at $\theta_{\star}\in\Theta$ if $F_{\theta}\neq F_{\theta_{\star}}$ for all $\theta\neq\theta_{\star}$.
Finally, as noted in the introduction, we assume that the Euclidean and Hellinger metrics are equivalent on any parametric model under consideration. Accordingly, we will use the notions of Hellinger and Euclidean consistency interchangeably throughout the paper, without further mention. In particular, Hellinger consistency is to be understood as in (1), with the Euclidean metric replaced by $d_{h}$ in the definition of the neighborhood $A_{\varepsilon}$.
3 Classical conditions for parametric posterior consistency
As discussed in Section 1, the predominant approach for establishing posterior consistency in parametric models is to verify that regularity conditions of the type proposed by Walker, (1969) are satisfied. In particular, restricting attention to the one-dimensional setting $\Theta\subseteq\mathbb{R}$, in addition to the basic requirements that $\Theta$ be closed and $\mathcal{F}_{\Theta}$ identifiable, Walker, (1969) required the following:
  1. W1.
    The set $\mathcal{O}_{\theta}:=\{x\in\mathbb{R}:f_{\theta}(x)=0\}$ is the same for all $\theta\in\Theta$.
  2. W2.
    For all $\theta\in\Theta$ and $x\in\mathbb{R}$, there exists a function $H_{\delta}(x,\theta)$ such that, for all $\delta>0$ small enough and all $\theta^{\prime}\in\mathbb{R}$ with $|\theta-\theta^{\prime}|<\delta$, the following conditions hold:
    $|\ln f_{\theta}(x)-\ln f_{\theta^{\prime}}(x)|$ $<H_{\delta}(x,\theta),$
    $\lim_{\delta\to 0}H_{\delta}(x,\theta)$ $=0,$
    $\lim_{\delta\to 0}\int_{\mathbb{R}}H_{\delta}(x,\theta)f_{\theta_{\star}}(x)\,\mathrm{d}x$ $=0,\quad\forall\theta_{\star}\in\Theta;$
  3. W3.
    If $\Theta$ is unbounded, then for all $\theta_{\star}\in\Theta$ and sufficiently large $M>0$, there exists a function $K_{M}(x,\theta_{\star})$ such that
    $\ln f_{\theta}(x)-\ln f_{\theta_{\star}}(x)<K_{M}(x,\theta_{\star})$
    for all $|\theta|>M$, with
    $\lim_{M\to\infty}\int_{\mathbb{R}}K_{M}(x,\theta_{\star})f_{\theta_{\star}}(x)\,\mathrm{d}x<0.$
Intuitively, assumption W1 ensures that the support of the model remains fixed across parameter values, thereby preventing singularities in likelihood ratios. Assumption W2 imposes a form of local uniform continuity on the log-likelihood function with respect to the parameter, captured via a modulus of continuity that vanishes both pointwise and in expectation under any true model. Finally, assumption W3 provides control over the relative tail behavior of the log-likelihood, requiring that log-likelihood ratios decay sufficiently fast—in an integrable manner—as the parameter value diverges.
Walker, (1969) showed that assumptions W1–W3, together with a positive and continuous prior density, imply posterior consistency at any $\theta_{\star}\in\Theta$ by ensuring MLE consistency and controlling the likelihood ratio away from the true parameter value. While technically valid—and while the positivity of the prior density may, under mild regularity conditions, be viewed as equivalent to the KL support condition of Schwartz, (1964)—these assumptions are quite specific and restrictive (see, e.g., Section 5 for a demonstration of this point). More importantly, their formulation is not informed by any clear connection to the core mechanisms that govern posterior consistency. Before presenting our own alternative treatment, we briefly review another classical approach to parametric consistency which, although potentially more broadly applicable, shares many of the same underlying limitations.
3.1 An alternative MLE-based strategy
Walker and Hjort, (2001) approached the problem of posterior consistency as follows. Taking $A_{\varepsilon}$ to be a Hellinger $\varepsilon$-ball around $f_{\theta_{\star}}$, consider
$\Pi(A_{\varepsilon}^{c}\mid X_{1:n})=\frac{\int_{A_{\varepsilon}^{c}}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)}{\int_{\Theta}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)}$
for sufficiently small $\varepsilon>0$. Standard results (see, e.g., Barron et al.,, 1999, Lemma 4) provide suitable lower bounds for the denominator as long as $\theta_{\star}\in\textnormal{KLS}(\Pi)$. For the numerator, letting $\hat{\theta}_{n}$ denote an MLE, one can write
$\int_{A_{\varepsilon}^{c}}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)\leq\prod_{i=1}^{n}\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\int_{A_{\varepsilon}^{c}}\prod_{i=1}^{n}\left(\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\Pi(\mathrm{d}\theta).$
While the second factor is readily shown to decay exponentially in $n$, the key idea in Walker and Hjort, (2001) is to require that
$\prod_{i=1}^{n}\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}<e^{cn/2}\quad\iff\quad\frac{1}{n}\sum_{i=1}^{n}\ln\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)<c$ (2)
eventually a.s.-$F_{\theta_{\star}}^{\infty}$ for all $c>0$, which arises if
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\ln\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)=0$
a.s.-$F_{\theta_{\star}}^{\infty}$. Under this condition, one obtains $\lim_{n\to\infty}\Pi(A_{\varepsilon}^{c}\mid X_{1:n})=0$ a.s.-$F_{\theta_{\star}}^{\infty}$ for all $\varepsilon>0$, thus establishing posterior consistency.
The condition in (2) is often satisfied in regular parametric models, for instance as guaranteed by the classical theorem of Wilks, (1938). Nonetheless, this proof technique inherits the same core limitations as that of Walker, (1969): it relies on potentially stringent regularity conditions to guarantee MLE behavior (Kiefer and Wolfowitz,, 1956; van de Geer,, 2000; van der Vaart,, 2000), and achieves consistency through a sequence of technical bounds, rather than by appealing to a conceptual framework that directly explains why the posterior should concentrate. In the next section, we move beyond these approaches and develop a more principled and flexible route to consistency, rooted in the concept of sequential identifiability.
4 Sequential identifiability, oscillations and posterior consistency
As mentioned in the introduction, our starting point is the groundbreaking result of Schwartz, (1964). Given its central role, we state it here (without proof).
Theorem 1 (Schwartz,, 1964).
Let $\theta_{\star}\in\textnormal{KLS}(\Pi)$. Then the posterior is weakly consistent at $\theta_{\star}$, that is,
$\lim_{n\to\infty}\Pi(\{\theta\in\Theta:d_{w}(F_{\theta},F_{\theta_{\star}})>\varepsilon\}\mid X_{1:n})=0$
a.s.-$F_{\theta_{\star}}^{\infty}$ for all $\varepsilon>0$.
While remarkable, Theorem 1 is not sufficient to address consistency in density estimation, as the topology induced by $d_{w}$ is too weak to meaningfully capture closeness between densities. In particular, it is possible to construct sequences of densities that do not converge in the Hellinger sense, yet whose associated cdfs converge pointwise to that of a continuous random variable—i.e., they converge weakly.
This phenomenon, which will be analyzed in greater detail in Theorem 3 and illustrated through a concrete model in Section 5, arises from the fact that a sequence of densities may oscillate indefinitely in a manner that precludes convergence in Hellinger distance, while still approximating the target cdf in distribution. In such cases, although the limit in density does not exist, the sequence may nonetheless approximate the cdf of the density around which it oscillates.
To address this potential gap between the two notions of convergence, a natural strategy is to rule out the possibility of such pathological approximations occurring within the parametric family of interest. The next definition, fundamental to our analysis, formalizes this idea.
Definition 3.
[…]
Intuitively, the concept of sequential identifiability implies that there exists no region of the Euclidean parameter space where $F_{\theta_{\star}}$ can be sequentially approximated in distribution, except in arbitrarily small neighborhoods of $\theta_{\star}$ itself. For example, if $\Theta=[0,\infty)$, sequential identifiability precludes the possibility that $F_{\theta}$ approximates $F_{\theta_{\star}}$ in distribution as $\theta\to\infty$. Moreover, observe that if $(\theta_{j})_{j\in\mathbb{N}}$ is a sequence such that $\lim_{j\to\infty}\theta_{j}=\theta\neq\theta^{\prime}$, then our assumption that Euclidean convergence implies Hellinger convergence (and therefore weak convergence) ensures that, under sequential identifiability on all of $\Theta$, we must have $F_{\theta}\neq F_{\theta^{\prime}}$. Thus, sequential identifiability may be interpreted as a strengthening of the standard notion of identifiability (Definition 2), requiring that no distribution in the model appears at multiple locations of the parameter space, once the latter is suitably extended to include weak limits.
The next theorem is the central result of our analysis, establishing a key connection between sequential identifiability and posterior consistency.
Theorem 2.
If $\theta_{\star}\in\textnormal{KLS}(\Pi)$ and $\mathcal{F}_{\Theta}$ is sequentially identifiable at $\theta_{\star}$, the posterior is consistent at $\theta_{\star}$. In particular, the posterior predictive density
$\hat{f}_{n}:=\int_{\Theta}f_{\theta}\,\Pi(\mathrm{d}\theta\mid X_{1:n})$
is a consistent estimator of $f_{\theta_{\star}}$.
Theorem 2 is noteworthy in that it establishes posterior consistency under very mild assumptions: sequential identifiability of the model and a prior that assigns positive mass to any KL neighborhood of the true parameter, the latter being a standard well-specification condition. The intuition behind the derivation of Theorem 2 is as follows: the KL support condition, via Schwartz’s theorem, ensures that the posterior concentrates in weak neighborhoods of $F_{\theta_{\star}}$. Sequential identifiability then rules out the possibility that such concentration occurs around points in the parameter space other than that corresponding to the true density, thereby yielding posterior consistency.
A practical implication of this result is that, to establish parameter consistency at some $\theta_{\star}$, it suffices to verify that the model space—augmented by its weak limit points—does not contain $F_{\theta_{\star}}$ at more than one location (e.g., both within the parameter space and along a sequence $(\theta_{j})_{j\in\mathbb{N}}$ with $\|\theta_{j}\|\to\infty$). As we show in the next subsection, the possibility of such a scenario is tied to the presence of pathological oscillations of the model densities around $f_{\theta_{\star}}$, which is highly unlikely in any parametric modeling setting unless the modeler possesses specific knowledge of $f_{\theta_{\star}}$ and deliberately constructs the model to fail sequential identifiability at $\theta_{\star}$.
Another insight from Theorem 2 arises from the following observations. If the posterior fails to be consistent at some $\theta_{\star}\in\Theta$, Theorem 2 implies the existence of a region in the augmented parameter space that induces sequential unidentifiability at $\theta_{\star}$, with the posterior accumulating with positive probability around such a weak limit point, separated from $\theta_{\star}$ in the Euclidean sense. Under the basic assumption of identifiability, however, inconsistency cannot result in posterior concentration around a different point within the parameter space, as this would contradict identifiability. Rather, the posterior mass must shift toward a region in the augmented space, lying outside the original parameter space, where the true distribution is still recovered. In other words, the posterior does not simply “miss” the true parameter by concentrating around a nearby but incorrect value; instead, it shifts toward regions in the augmented space proper, such as points at infinity, which correspond to weak limits of the true distribution. This observation suggests a practical heuristic for diagnosing strong consistency: in addition to verifying sequential identifiability, one may examine whether the posterior remains confined within the parameter space, rather than drifting toward, for instance, infinity. In light of our discussion, such stability would provide evidence in support of consistency.
4.1 The role of oscillations
While sequential identifiability is the central concept of our theoretical analysis, it is crucial to understand the implications of its failure. The next theorem addresses this question by examining the behavior of a sequence of density functions that do not converge in the Hellinger metric (which, in our parametric setting, is assumed to be equivalent to non-convergence of the associated parameter sequence), yet whose corresponding distribution functions do converge to a proper limit.
To state the next result, we first establish the following pieces of terminology. For any two densities $f$ and $g$ such that $A:=\{x\in\mathbb{R}:g(x)>f(x)\}$ is open, we say that $g$ oscillates $O$ times around $f$ if $O\in\mathbb{N}$ is the minimum number of disjoint intervals $(a_{1},b_{1}),\dots,(a_{O},b_{O})$ such that we can write
$A=\bigcup_{i=1}^{O}(a_{i},b_{i}).$
Notice that the openness of $A$ implies that there exists a decomposition of it into countably many disjoint open intervals, and we call the above expression the minimal decomposition of $A$.
Theorem 3.
Let $g,f_{1},f_{2},\dots$ be densities such that the sets
$A_{j}:=\{x\in\mathbb{R}:f_{j}(x)>g(x)\},\quad B_{j}:=\{x\in\mathbb{R}:g(x)>f_{j}(x)\}$
are open for all $j\in\mathbb{N}$ [^2] While openness of $A_{j}$ and $B_{j}$ is a minimal requirement needed in the proof of the result, notice that any set of densities $g,f_{1},f_{2},...$ that are continuous on a common support satisfies the assumption. Moreover, assume that (i) $d_{w}(F_{j},G)\to 0$ as $j\to\infty$, and (ii) $d_{h}(f_{j},g)\geq\varepsilon>0$ for all $j\in\mathbb{N}$. Then the number of oscillations $O_{j}$ of $f_{j}$ around $g$ tends to infinity as $j\to\infty$.
The preceding result implies that, if sequential identifiability fails at some $\theta\in\Theta$, then the model must contain a sequence of densities that oscillate arbitrarily frequently around $f_{\theta}$. Figure 1 provides a graphical illustration of how such a sequence would behave. Consequently, ruling out this type of pathological behavior is sufficient to guarantee posterior consistency. In practice, posterior inconsistency due to such behavior could only arise if the modeler possessed precise knowledge of the true density and deliberately introduced carefully designed oscillations around it. We observe that this is an implausible scenario in any reasonable parametric modeling context, which makes our conditions for parametric consistency virtually universal.
The role of oscillations will be further explored in Section 5 using a simple illustrative parametric model. Before doing so, we review a few common parametric families and show how consistency can be easily established using the conditions introduced in this section.
Figure 1: The initial segment of a sequence $f_{1},f_{2},\dots$ of indefinitely oscillating densities (solid gray, top row), whose corresponding sequence of cdfs (solid gray, bottom row) converges to a proper limiting distribution (dashed black). Such convergence is only possible because the density sequence exhibits increasingly frequent oscillations around the density associated with the limiting cdf.
4.2 Example 1: exponential families
Assume that, for each $\theta\in\Theta\subseteq\mathbb{R}^{p}$, the associated density satisfies
$f_{\theta}(x)\propto h(x)\exp\left\{\theta^{\top}T(x)\right\},\quad x\in\mathbb{R}.$
Then $\mathcal{F}_{\Theta}$ defines a $d$-dimensional exponential family with sufficient statistics $T(x)\in\mathbb{R}^{p}$ (Efron,, 2022). This general form includes several classical parametric models for continuous data—such as the Gaussian, exponential, gamma, beta, Laplace, Rayleigh, Weibull, and von Mises distributions—for which sequential identifiability (and therefore posterior consistency) is readily verified.
4.3 Example 2: uniform distribution
[…]
4.4 Example 3: finite mixture models
Consider a normal mixture model with a finite number $K$ of components. For ease of exposition, we circumvent the usual parameter identifiability issues associated with mixtures (Teicher,, 1963) and work directly on the space of mixture densities equipped with the Hellinger metric: [^3] Nevertheless, by imposing standard identifiability constraints on the mixture parameters, the following analysis can be extended to the usual Euclidean setting. We also note that the same line of reasoning applies to mixtures with more general kernel families. specifically, the $K$-component normal mixture model is defined as
$\mathcal{F}^{K}:=\left\{\sum_{k=1}^{K}w_{k}\,\lambda_{k}^{1/2}\phi\left(\lambda_{k}^{1/2}(\cdot-\mu_{k})\right):(\bm{w},\bm{\mu},\bm{\lambda})\in\Delta_{K}\times\mathbb{R}^{K}\times(0,\infty)^{K}\right\},$
where $\Delta_{K}$ denotes the $(K-1)$-dimensional simplex and $\phi(x):=(2\pi)^{-1/2}e^{-x^{2}/2}$ for all $x\in\mathbb{R}$. Since any two densities $f,g\in\mathcal{F}^{K}$ are continuous, we may invoke Theorem 3 to establish sequential identifiability, and hence posterior consistency. Indeed, for any fixed $K\in\mathbb{N}$, the number of oscillations of $f$ around $g$ is bounded above by a finite constant (common to all $f,g\in\mathcal{F}^{K}$), implying that weak convergence within $\mathcal{F}_{\Theta}$ entails Hellinger convergence. Consequently, by Theorem 2, any density in the KL support of the prior will exhibit posterior consistency.
5 An illustrative model
We are now in a position to illustrate our theory with a simple yet instructive parametric model. Specifically, we restrict the sample space to $[0,1]$ and consider the one-dimensional parameter space $\Theta=[0,\infty)$, where the family $\mathcal{F}_{\Theta}$ consists of densities of the form
$f_{\theta}(x)=\frac{1+\cos(\theta x)}{1+\sin(\theta)/\theta}.$ (3)
To be precise, we adopt the convention that $f_{0}$ is defined as the continuous extension of the above expression as $\theta\to 0$, yielding $f_{0}(x)=1$ for all $x\in[0,1]$—that is, $F_{0}$ corresponds to the uniform distribution. The resulting family of cdfs is of the form
$F_{\theta}(x)=\frac{x+\sin(\theta x)/\theta}{1+\sin(\theta)/\theta}.$
As illustrated in Figure 2, increasing values of $\theta$ induce more pronounced oscillations in the density around the value 1, while the corresponding cdf converges to $F_{0}$ (cf. Figure 1 in Section 4). We summarize these and other basic properties of the model in the next proposition.
Figure 2: Density functions (solid gray, top row) and cdfs (solid gray, bottom row) in the illustrative parametric model, shown for increasing values of the parameter $\theta$. As $\theta$ grows, the density becomes increasingly oscillatory while the cdf converges to $F_{0}$ (dashed black).
Proposition 1.
The parametric family $\mathcal{F}_{\Theta}$ defined by (3) satisfies the following properties:
  1. 1.
    The Euclidean and Hellinger metrics are equivalent on $\mathcal{F}_{\Theta}$;
  2. 2.
    For any $\theta\geq 0$ and prior $\Pi$ on $\Theta$, if $\Pi(A)>0$ for every neighborhood $A$ of $\theta$, then $\theta\in\textnormal{KLS}(\Pi)$;
  3. 3.
    $\mathcal{F}_{\Theta}$ is sequentially identifiable at all $\theta>0$;
  4. 4.
[…]
The first property confirms that the assumption of equivalence between the Euclidean and Hellinger metrics, made throughout this article, holds in this setting as well. This allows us to move freely between the two notions of convergence without loss of generality. The second property implies that, for the KL support condition to hold at all parameter values, it suffices for the prior to have full support on $\Theta$—a very mild requirement. Combined with the third property, this enables us to invoke Theorem 2 and obtain the following corollary.
Corollary 1.
If the prior $\Pi$ has full support on $[0,\infty)$, then the posterior is consistent at all $\theta>0$.
This result is remarkable because, as we demonstrate in Subsection 5.1 below, the family $\mathcal{F}_{\Theta}$ fails to satisfy classical sufficient conditions for consistency proposed in earlier literature (cf. Section 3). Nonetheless, our simple criterion of sequential identifiability seamlessly yields posterior consistency at all $\theta>0$.
Due to the fourth property in Proposition 1, however, the same argument does not extend to $\theta=0$, as the model is sequentially unidentifiable at that point. In particular, while a prior with full support guarantees posterior concentration in weak neighborhoods of $F_{0}$ (thanks to Schwartz’s theorem), such concentration may occur in the “wrong” region of the parameter space—namely, toward infinity. In line with Theorem 3, this happens because this cosine-based model oscillates around the uniform density with increasing frequency as the parameter $\theta$ diverges. It is worth emphasizing the specificity of the model’s construction in producing such behavior: the oscillations are finely tuned to occur around the uniform density and do not induce inconsistency at any other parameter value. This further reinforces our earlier argument that, although violations of sequential identifiability are theoretically possible, they are highly unlikely to arise in any reasonably specified parametric model, let alone pose a threat to posterior consistency. Indeed, sequential unidentifiability at the true parameter value would require the modeler to possess precise knowledge of the data-generating process and to precisely engineer the statistical model to behave pathologically around it—as we have done in this example.
5.1 Failure of classical conditions
While our principled analysis immediately established consistency for all $\theta>0$, we now show that the oscillatory behavior of the densities defined in (3) violates the classical regularity conditions commonly used to ensure posterior consistency. As a consequence, these existing results are inadequate for analyzing posterior convergence in this model.
In particular, with respect to the conditions introduced by Walker, (1969), we have the following proposition.
Proposition 2.
The parametric model in (3) fails assumptions W1 and W3 (cf. Section 3).
Heuristically, the failure of assumption W1 stems from the model’s oscillations causing the density to vanish at varying points on the support. The failure of assumption W3, on the other hand, arises from the persistence of oscillations in the likelihood function even as $\theta\to\infty$.
More fundamentally, we next establish that, under this model, the MLE is inconsistent at all data-generating parameter values. This result reinforces the limitations of the approach developed by Walker, (1969), which relies on model regularity to infer posterior consistency from MLE consistency. The inconsistency of the MLE renders the methodology of Walker and Hjort, (2001) inapplicable as well, since that framework also presupposes MLE consistency.
Theorem 4.
For all true data-generating parameters $\theta_{\star}\geq 0$, maximum likelihood estimation is inconsistent. Specifically, for all $M>0$,
$\max_{\theta\in[0,M]}\prod_{i=1}^{n}f_{\theta}(X_{i})<\sup_{\theta\geq 0}\prod_{i=1}^{n}f_{\theta}(X_{i})$
ultimately a.s.-$F_{\theta_{\star}}^{\infty}$.
The proof of Theorem 4 relies on a classical number-theoretic result, namely Dirichlet’s simultaneous approximation theorem (Schmidt,, 1980). Heuristically, the argument proceeds in two steps. First, we show that any MLE based on $n$ observations must achieve a product likelihood of at least $2^{n}$, as it is always possible to select a value of $\theta$ such that each data point aligns with a peak of the cosine oscillations (which, as can be readily verified, reach values of around 2). Second, we demonstrate that, for all sufficiently large $n$, with probability 1, no likelihood maximizer restricted to $[0,M]$ for any fixed $M>0$ can attain this lower bound. This establishes the asymptotic failure of the maximum likelihood principle—and, consequently, of classical approaches to posterior consistency—at all $\theta_{\star}\geq 0$.
5.2 Consistency at the uniform density
[…]
The first approach we consider is the one introduced in the seminal papers by Ghosal et al., (1999) and Barron et al., (1999), which establishes exponential decay of the posterior mass outside Hellinger neighborhoods of $\theta_{\star}$ by splitting the complement of such neighborhoods into two $n$-dependent regions: the first has slowly growing complexity—e.g., measured in terms of Hellinger metric or upper-bracketing entropy (van de Geer,, 2000; Wainwright,, 2019)—forcing the likelihood ratio to decay exponentially; the second region is directly assumed to have exponentially small prior (hence posterior) mass. Together with the KL support assumption and an upper-bound of $d_{h}(f_{\theta},f_{\theta^{\prime}})$ by the corresponding Euclidean distance (see the supplementary material for more details), this leads to the following result.
Proposition 3.
Assume that $0\in\textnormal{KLS}(\Pi)$ and that, for some function [^4] Valid choices of $\varphi$ include $\varphi(t)=t^{1+\beta}$ and $\varphi(t)=\beta\exp(t)$ for some $\beta>0$. The latter condition, for example, is satisfied by an exponential prior. $\varphi:(0,\infty)\to(0,\infty)$ such that $\lim_{t\to\infty}\varphi(t)/t=\infty$, the prior cdf satisfies
$1-\Pi(\theta)\leq e^{-\varphi(\ln\theta)}$
for all sufficiently large $\theta>0$. Then the posterior is consistent at $\theta_{\star}=0$.
A second approach is that of Walker, (2004), who showed that posterior consistency holds as long as the parameter space can be covered by sets whose prior masses satisfy a suitable summability condition. In our setting, this yields the following result.
Proposition 4.
Assume that $0\in\textnormal{KLS}(\Pi)$ and that the prior density $\pi$, for all sufficiently large $\theta>0$, is decreasing and satisfies
$\pi(\theta)\lesssim\frac{1}{\theta^{2}(\ln\theta)^{2+\beta}}$ (4)
for some $\beta>0$. Then the posterior is consistent at $\theta_{\star}=0$.
Finally, we consider a third strategy that leverages insights from our analysis of oscillations in the product likelihood, as well as structural properties of the model, to establish consistency under even weaker assumptions on the prior. Specifically, we exploit the observation that the product likelihood, when restricted to values of $\theta$ smaller than $e^{cn}$ (for some $c>0$), is asymptotically well behaved: in this region, oscillations eventually vanish, and the restricted MLE exhibits regular behavior. Instead, for the region where $\theta>e^{cn}$, any prior with sub-polynomial tails suffices to ensure the necessary posterior mass decay, leading to the following result.
Theorem 5.
Assume that $0\in\textnormal{KLS}(\Pi)$ and that the prior density $\pi$, for all sufficiently large $\theta>0$, satisfies
$\pi(\theta)\lesssim\theta^{-(1+\alpha)}$
for some $\alpha>0$. Then the posterior is consistent at $\theta_{\star}=0$.
6 Discussion
In this work, we introduced a general framework for studying posterior consistency in parametric models, based on the simple yet powerful notion of sequential identifiability. Departing from the regularity-based assumptions commonly found in the literature, our approach builds on the observation that, by the seminal result of Schwartz, (1964), weak consistency is guaranteed under a mild KL support condition. Sequential identifiability then emerges as a minimal additional assumption that enables the translation of weak consistency into parameter consistency. We further showed that inconsistency can arise only in the presence of oscillatory behavior in the model—an insight that, together with the sequential identifiability condition itself, leads to consistency results under assumptions significantly weaker than those required by classical theory. This perspective enables the study of models beyond the reach of existing approaches, including the parametric family we constructed and analyzed in detail.
Our work has a multiplicity of ramifications. Most directly, it offers novel and accessible guidance for the applied Bayesian statistician seeking to understand the asymptotic behavior of parametric models of interest. Building on our analysis of sequential identifiability, our message to the modeler is straightforward: rather than verifying a list of regularity conditions, simply avoid introducing arbitrarily oscillating densities into the model in the first place. If, for some reason, such oscillations are unavoidable, there is still little cause for concern about inconsistency—unless one possesses precise knowledge of the true data-generating distribution and is willing to meticulously design pathological oscillations centered around it. However, in such a scenario the need for statistical inference itself becomes questionable: after all, if the modeler already knows so much about the true distribution as to craft these adversarial features, inference would hardly be necessary. As we have illustrated, our carefully constructed example required exactly this kind of contrived design to challenge consistency, yet it remained sequentially identifiable at all but a single parameter value. Moreover, even at the parameter value where sequential identifiability fails, we showed that posterior consistency can still be salvaged through a fine-grained analysis of the model’s oscillatory behavior and suitable control via the prior.
Second, our analysis bears strong connections with the literature on nonparametric posterior consistency. Although not explicitly, the sieve complexity conditions introduced by Barron et al., (1999), Ghosal et al., (1999), and Walker, (2004) aim to address a similar underlying issue: in those works, oscillatory behavior in the density model—leading to sequential unidentifiability—is quantified and controlled, for instance, via the Hellinger metric entropy of sets where most of the prior mass is concentrated. However, it is worth emphasizing that, in nonparametric models, sequential unidentifiability is almost always inherent. For example, infinite Gaussian mixtures can weakly approximate any distribution on the real line while failing to converge in density, due to components with rapidly increasing precision parameters. In such settings, sequential identifiability is not a reasonable assumption, and a more refined analysis of the interaction between the model’s oscillatory behavior and the prior is necessary.
Finally, our analysis—particularly the construction of a model based on cosine oscillations—sheds new light on a well-known counterexample by Barron et al., (1999), in which the posterior is weakly consistent but not consistent in Hellinger distance. In that example, positive prior mass is placed on each member of a family of discontinuous, oscillatory densities on $[0,1]$ that alternate between values of 0 and 2 across disjoint intervals, thereby weakly approximating the uniform density. In our terminology, this corresponds to a model that is sequentially unidentifiable at the uniform density, and the mechanism driving inconsistency is closely related to our example, in which continuous oscillations occur between 0 and (approximately) 2 around the uniform density. Nevertheless, our model has been shown to achieve consistency at the uniform distribution under mild tail conditions on the prior, while the carefully placed prior mass on discontinuous, oscillatory densities in the Barron et al., (1999) example results in inconsistency. Although beyond the scope of this study, these observations suggest the possibility that additional mild assumptions on the nature of oscillations (e.g., continuity) or on the prior (e.g., no positive mass on single parameters or densities) may help to rule out inconsistent posterior behavior even in nonparametric settings.
Supplementary Material for “Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability"
In this supplementary material, we present the proofs of all the theoretical results from the main text of the article.
Appendix A Proof of Theorem 2
By Theorem 1, the KL support assumption implies that, for all $\varepsilon>0$,
$\Pi(\{\theta\in\Theta:d_{w}(F_{\theta_{\star}},F_{\theta})\leq\varepsilon\}\mid X_{1:n})\to 1$
a.s.-$F_{\theta_{\star}}^{\infty}$. Specifically, choosing any positive decreasing sequence $\varepsilon_{j}\to 0$,
$F_{\theta_{\star}}^{\infty}\left(x_{1:\infty}\in\mathbb{R}^{\mathbb{N}}:\,\Pi(\{\theta\in\Theta:d_{w}(F_{\theta_{\star}},F_{\theta})\leq\varepsilon_{j}\}\mid x_{1:n})>1-\delta\textnormal{ ultimately}\right)=1.$
for all $\delta>0$ and $j\in\mathbb{N}$. Now fix $j\in\mathbb{N}$ and assume per contra that the posterior is not consistent in the Euclidean metric, so that there exist $\varepsilon>0$ and $\delta>0$ such that
[…]
Because $F_{\theta_{\star}}^{\infty}$ is a probability measure, the two above expressions imply that there exists some $x_{1:\infty}$ such that
$\Pi(\{\theta\in\Theta:d_{w}(F_{\theta_{\star}},F_{\theta})\leq\varepsilon_{j}\}\mid x_{1:n})$ $>1-\delta\quad\textnormal{ultimately},$
[…]
This implies the existence of some $n\in\mathbb{N}$ such that
$\Pi(\{\theta\in\Theta:d_{w}(F_{\theta_{\star}},F_{\theta})\leq\varepsilon_{j}\}\mid x_{1:n})$ $>1-\delta,$
$\Pi(\{\theta\in\Theta:\|\theta_{\star}-\theta\|>\varepsilon\}\mid x_{1:n})$ $>\delta.$
Because $\Pi(\cdot\mid x_{1:n})$ is a probability measure, the last two inequalities imply the existence of some $\theta_{j}\in\Theta$ such that both $d_{w}(F_{\theta_{\star}},F_{\theta_{j}})\leq\varepsilon_{j}$ and $\|\theta_{\star}-\theta_{j}\|>\varepsilon$. Because we can repeat the above argument for all $j\in\mathbb{N}$, we can construct a sequence $(\theta_{j})_{j\in\mathbb{N}}$ such that $d_{w}(F_{\theta_{\star}},F_{\theta_{j}})\leq\varepsilon_{j}\to 0$ but $\inf_{j\in\mathbb{N}}\|\theta_{\star}-\theta_{j}\|\geq\varepsilon>0$. This leads to a contradiction of sequential identifiability at $\theta_{\star}$, concluding the proof of posterior consistency.
As for the consistency of $\hat{f}_{n}$, Jensen’s inequality applied to the convex map $f\mapsto d_{h}(f,f_{\theta_{\star}})$, together with the definition $A_{\varepsilon}:=\{\theta\in\Theta:d_{h}(f_{\theta},f_{\theta_{\star}})<\varepsilon\}$ for $\varepsilon>0$, implies
$d_{h}(\hat{f}_{n},f_{\theta_{\star}})$ $\leq\int_{\Theta}d_{h}(f_{\theta},f_{\theta_{\star}})\,\Pi(\mathrm{d}\theta\mid X_{1:n})$
$=\int_{A_{\varepsilon}}d_{h}(f_{\theta},f_{\theta_{\star}})\,\Pi(\mathrm{d}\theta\mid X_{1:n})+\int_{A_{\varepsilon}^{c}}d_{h}(f_{\theta},f_{\theta_{\star}})\,\Pi(\mathrm{d}\theta\mid X_{1:n})$
$\leq\varepsilon+\sqrt{2}\,\Pi(A_{\varepsilon}^{c}\mid X_{1:n}).$
The second term converges to zero a.s-$F_{\theta_{\star}}^{\infty}$ for any $\varepsilon>0$, by posterior consistency, while the first term can be made arbitrarily small. This completes the proof.
Appendix B Proof of Theorem 3
For convenience, recall the definition of the Lévy-Prokhorov distance between distributions $F_{j}$ and $G$:
$d_{w}(F_{j},G):=\inf\left\{\delta>0:\,\forall A\in\mathscr{B}(\mathbb{R}),\,F_{j}(A)\leq G(A^{\delta})+\delta,\,G(A)\leq F_{j}(A^{\delta})+\delta\right\},$
where $A^{\delta}:=\{x\in\mathbb{R}:\exists y\in A,\,|x-y|<\delta\}$. Therefore, calling $\delta_{j}:=d_{w}(F_{j},G)$, by assumption we get [^5] Notice that, if the infimum in the definition of $d_{w}$ is not attained by $\delta_{j}$, it suffices to replace $\delta_{j}$ with $\delta_{j}+\varepsilon_{j}$, for some non-negative sequence $\varepsilon_{j}\to 0$, in the following analysis. Therefore, without loss of generality, we continue to work with $\delta_{j}$.
$F_{j}(A)\leq G(A^{\delta_{j}})+\delta_{j}\textnormal{ and }G(A)\leq F_{j}(A^{\delta_{j}})+\delta_{j},\qquad\forall A\in\mathscr{B}(\mathbb{R}).$ (5)
Now recall that, for probability measures on $\mathbb{R}$ admitting a density, the Hellinger distance is topologically equivalent to the total variation distance $d_{TV}$, so assume without loss of generality that
$\varepsilon\leq d_{TV}(F_{j},G):=\sup_{A\in\mathscr{B}(\mathbb{R})}|F_{j}(A)-G(A)|\equiv\frac{1}{2}\int_{\mathbb{R}}|f_{j}(x)-g(x)|\mathrm{d}x.$
The above characterization of $d_{TV}$ implies that the $\sup$ in its definition is attained either by $A_{j}$ at $F_{j}(A_{j})-G(A_{j})\geq\varepsilon$, or by $B_{j}$ at $G(B_{j})-F_{j}(B_{j})\equiv F_{j}(A_{j})-G(A_{j})\geq\varepsilon$. Without loss of generality, assume that the $\sup$ is attained by $A_{j}$ and
$F_{j}(A_{j})-G(A_{j})\geq\varepsilon,$ (6)
the other case being perfectly symmetric. Denote by
$A_{j}=\bigcup_{i=1}^{O_{j}}(a_{ij},b_{ij}),$
the minimal decomposition of $A_{j}$, which exists by the assumed openness of $A_{j}$. Then Equations (5) and (6) combine into $\varepsilon\leq G(A_{j}^{\delta_{j}}\setminus A_{j})+\delta_{j}$, where
$A_{j}^{\delta_{j}}\setminus A_{j}\subseteq\bigcup_{i=1}^{O_{j}}\Big{(}[a_{ij}-\delta_{j},a_{ij}]\cup[b_{ij},b_{ij}+\delta_{j}]\Big{)}$
is such that
$G(A_{j}^{\delta_{j}}\setminus A_{j})\leq 2O_{j}\sup_{x\in\mathbb{R}}\left\{G(x+\delta_{j})-G(x)\right\}.$
Therefore
$\varepsilon\leq 2O_{j}\sup_{x\in\mathbb{R}}\left\{G(x+\delta_{j})-G(x)\right\}+\delta_{j}$
for all $j\in\mathbb{N}$. Because the distribution $G$ is a continuous, bounded and monotonically increasing function, it is also uniformly continuous, so $\lim_{j\to\infty}\delta_{j}=0$ implies $\lim_{j\to\infty}\sup_{x\in\mathbb{R}}\left\{G(x+\delta_{j})-G(x)\right\}=0$. This, together with the last expression, yields $\lim_{j\to\infty}O_{j}=\infty$.
Appendix C Proof of Proposition 1
We prove that each property holds separately.
C.1 Proof of property 1
To prove the equivalence of the Euclidean and Hellinger metrics, we rely on the following lemma.
Lemma 1.
For all $\theta,\theta^{\prime}\geq 0$, $d_{h}(f_{\theta},f_{\theta^{\prime}})\leq\min\left\{\sqrt{2},\,|\theta-\theta^{\prime}|\right\}$.
Proof.
$d_{h}(f_{\theta},f_{\theta^{\prime}})\leq\sqrt{2}$ holds by design, while checking the other upper-bound is straightforward once one verifies that the family of functions $\{\theta\mapsto\sqrt{f_{\theta}(x)}:x\in[0,1]\}$ is uniformly 1-Lipschitz continuous. To get this, it suffices to check that
$\left|\frac{\partial}{\partial\theta}\sqrt{f_{\theta}(x)}\right|=\left|\frac{x\sin(\theta x)}{2\sqrt{1+\cos(\theta x)}\sqrt{1+\sin(\theta)/\theta}}+\frac{\sqrt{1+\cos(\theta x)}(\theta\cos(\theta)-\sin(\theta))}{2\theta^{2}(1+\sin(\theta)/\theta)^{3/2}}\right|,$
which exists almost everywhere for all $x\in[0,1]$, is bounded above by 1 for all $x\in[0,1]$ and all $\theta\geq 0$ at which it exists, and moreover that $\theta\mapsto\sqrt{f_{\theta}(x)}$ is absolutely continuous on $[0,\infty)$ for all $x\in[0,1]$. Plugging this into the definition of $d_{h}$ yields the desired result. ∎
In particular, the preceding lemma shows that $d_{h}(f_{\theta},f_{\theta^{\prime}})\to 0$ if $|\theta-\theta^{\prime}|\to 0$. The reverse implication follows from the fact that, for $\theta>0$, Hellinger convergence to $f_{\theta}$ implies weak convergence to $F_{\theta}(x)=(x+\sin(\theta x)/\theta)/(1+\sin(\theta)/\theta)$, which can only happen if the parameter value converges to $\theta$ itself. As for $\theta=0$, instead, property 4 shows that weak convergence to $F_{0}$ can only happen as parameter values converge to 0 or as they diverge to $\infty$. However, in the latter case, Hellinger convergence fails (see again property 4), proving that $d_{h}(f_{\theta},f_{0})\to 0$ implies $\theta\to 0$.
C.2 Proof of property 2
Our aim is to show that $\lim_{\theta^{\prime}\to\theta}\textnormal{KL}(f_{\theta},f_{\theta^{\prime}})=0$ for all $\theta$. If that is the case, for all $\varepsilon>0$ small enough there exists $\delta>0$ so that $|\theta^{\prime}-\theta|<\delta$ implies $\textnormal{KL}(f_{\theta},f_{\theta^{\prime}})<\varepsilon$, or
$\{\theta^{\prime}\geq 0:|\theta^{\prime}-\theta|<\delta\}\subseteq\{\theta^{\prime}\geq 0:\textnormal{KL}(f_{\theta},f_{\theta^{\prime}})<\varepsilon\}.$
Because the smaller set has positive prior mass due to our assumption on the prior $\Pi$, one concludes that $\theta\in\textnormal{KLS}(\Pi)$.
So fix $\theta,\theta^{\prime}\geq 0$ and $x\in[0,1]$. Because $1+\sin(t)/t\in[\ell,2]$ for all $t\geq 0$ and some $\ell>0$, we have that
$\textnormal{KL}(f_{\theta},f_{\theta^{\prime}})$ $=\int_{0}^{1}\ln\left(\frac{f_{\theta}(x)}{f_{\theta^{\prime}}(x)}\right)f_{\theta}(x)\,\mathrm{d}x$
$\leq\ln\left(\frac{1+\sin(\theta^{\prime})/\theta^{\prime}}{1+\sin(\theta)/\theta}\right)+\frac{1}{\ell}\int_{0}^{1}\ln\left(\frac{1+\cos(\theta x)}{1+\cos(\theta^{\prime}x)}\right)(1+\cos(\theta x))\,\mathrm{d}x.$
By the continuity and boundedness away from 00 and $\infty$ of the function $\theta\mapsto 1+\sin(t)/t$, the first term converges to 00 as $\theta^{\prime}\to\theta$. As for the second addendum, notice that [^6] Here, we proceed under the assumption that $\theta>0$, the case $\theta=0$ being easily handled thanks to the strict positivity of densities with parameter lying in a neighborhood of 0.
$\int_{0}^{1}\ln\left(\frac{1+\cos(\theta x)}{1+\cos(\theta^{\prime}x)}\right)(1+\cos(\theta x))\,\mathrm{d}x=\frac{1}{\theta}\int_{0}^{\theta}\ln\left(\frac{1+\cos(s)}{1+\cos(\theta^{\prime}s/\theta)}\right)(1+\cos(s))\,\mathrm{d}s.$
Now fix $s\in[0,\theta]$ and obtain
$g(\theta^{\prime})$ $:=\ln\left(\frac{1+\cos(s)}{1+\cos(\theta^{\prime}s/\theta)}\right)(1+\cos(s))$
$=g(\theta)+g^{\prime}(\theta)(\theta^{\prime}-\theta)+\frac{g^{\prime\prime}(\phi)}{2}(\theta^{\prime}-\theta)^{2}$
$\leq g(\theta)+|g^{\prime}(\theta)||\theta^{\prime}-\theta|+\frac{|g^{\prime\prime}(\phi)|}{2}(\theta^{\prime}-\theta)^{2}$
by a second order Taylor expansion with remainder around $\theta^{\prime}=\theta$, where $\phi$ lies between $\theta$ and $\theta^{\prime}$. Clearly $g(\theta)=0$ and
$|g^{\prime}(\theta)|=\left|\frac{s\sin(\theta s/\theta)(1+\cos(s))}{\theta(1+\cos(\theta s/\theta))}\right|=\left|\frac{s\sin(s)}{\theta}\right|\leq 1.$
Moreover
$|g^{\prime\prime}(\phi)|$ $=\left|\frac{\cos(\phi s/\theta)}{1+\cos(\phi s/\theta)}+\left(\frac{\sin(\phi s/\theta)}{1+\cos(\phi s/\theta)}\right)^{2}\right|(1+\cos(s))\left(\frac{s}{\theta}\right)^{2}$
$\leq 2\left[\left|\frac{\cos(\phi s/\theta)}{1+\cos(\phi s/\theta)}\right|+\left(\frac{\sin(\phi s/\theta)}{1+\cos(\phi s/\theta)}\right)^{2}\right]$
where it is easily shown that, for $\phi$ sufficiently close to $\theta$ (hence $\theta^{\prime}$ sufficiently close to $\theta$), the two addenda in square brackets are upper-bounded by finite constants, uniformly over all $s\in[0,\theta]$. Therefore, we conclude that, for $\theta^{\prime}$ sufficiently close to $\theta$,
$\frac{1}{\theta}\int_{0}^{\theta}\ln\left(\frac{1+\cos(s)}{1+\cos(\theta^{\prime}s/\theta)}\right)(1+\cos(s))\,\mathrm{d}s$ $\lesssim|\theta-\theta^{\prime}|+(\theta-\theta^{\prime})^{2}\leq 2|\theta-\theta^{\prime}|$
where the constant in the inequality may depend on $\theta$. This completes the proof.
C.3 Proof of property 3
To prove sequential identifiability at all $\theta>0$, it is enough to observe that the model is identifiable and, as $\theta^{\prime}\to\infty$, $F_{\theta^{\prime}}$ does not converge weakly to $F_{\theta}$.
C.4 Proof of property 4
[…]
$\lim_{\theta\to\infty}F_{\theta}(x)=\lim_{\theta\to\infty}\frac{x+\sin(\theta x)/\theta}{1+\sin(\theta)/\theta}=x=F_{0}(x)$
for all $x\in[0,1]$. To show instead that there exists no Hellinger limit as $\theta\to\infty$, we proceed as follows. By what we just showed, if the Hellinger limit of $f_{\theta}$ as $\theta\to\infty$ existed, it would be $f_{0}(x)\equiv 1_{[0,1]}(x)$ (because Hellinger convergence implies weak convergence and weak limits are unique). That is, recalling that Hellinger and $L^{1}$ convergence are equivalent, we would have
$0=\lim_{\theta\to\infty}\int_{0}^{1}|f_{\theta}(x)-1|\mathrm{d}x=\int_{0}^{1}\lim_{\theta\to\infty}|f_{\theta}(x)-1|\mathrm{d}x,$
where the last equality comes from an application of the dominated convergence theorem to the bounded integrand $|f_{\theta}(x)-1|$. Therefore, $\lim_{\theta\to\infty}f_{\theta}(x)$ would exist for almost every $x\in[0,1]$, which we next show not to be the case. Let $\theta_{k}=2\pi k$ for all $k\in\mathbb{N}$ and fix $x\in[0,1]\setminus\mathbb{Q}$. Because $\sin(2\pi k)=0$ for all $k\in\mathbb{N}$, we can write $f_{\theta_{k}}(x)=1+\cos(2\pi kx)=1+\cos(2\pi\{kx\})$, where we denote by $\{r\}:=r-\lfloor r\rfloor$ the fractional part of $r\geq 0$. It is a well-known fact that, for irrational $x$, the set $\{\{kx\}:k\in\mathbb{N}\}$ is dense in $[0,1]$, and because continuous functions map dense sets to dense sets, $\{1+\cos(2\pi\{kx\}):k\in\mathbb{N}\}$ is dense in $[0,2]$. Therefore $\lim_{k\to\infty}f_{\theta_{k}}(x)$ does not exist for any $x\in[0,1]\setminus\mathbb{Q}$, which is a set of Lebesgue measure 1. This leads to a contradiction and concludes the proof.
Appendix D Proof of Proposition 2
Assumption W1 requires the set $\mathcal{O}_{\theta}:=\{x\in[0,1]:f_{\theta}(x)=0\}$ to be the same for all $\theta\geq 0$, which is clearly not true because
$\mathcal{O}_{\theta}=\{x\in[0,1]:\exists k\in\mathbb{N}_{0},\,\theta x=(2k+1)\pi\},$
which depends on $\theta$.
Assumption W3 requires that, for any $\theta_{\star}\geq 0$ and sufficiently large $M>0$, there exists a function $K_{M}(x,\theta_{\star})$ such that
$\ln f_{\theta}(x)-\ln f_{\theta_{\star}}(x)<K_{M}(x,\theta_{\star})$
for all $\theta>M$, with
$\lim_{M\to\infty}\int_{0}^{1}K_{M}(x,\theta_{\star})\,f_{\theta_{\star}}(x)\,\mathrm{d}x<0.$
From the proof of Theorem 4 (see the next section), we see that, for all $x\in[0,1]\setminus\mathcal{O}_{\theta_{\star}}$, $\delta>0$ and $M>0$, there exist infinitely many $\theta>M$ such that $\ln f_{\theta}(x)\geq\ln(2-\delta)$, so that any candidate $K_{M}(x,\theta_{\star})$ must satisfy
$K_{M}(x,\theta_{\star})\geq\ln(2-\delta)-\ln f_{\theta_{\star}}(x).$
From the proof of Lemma 2 below, it emerges that $\int_{0}^{1}f_{\theta_{\star}}^{2}(x)\mathrm{d}x<2$, so
$\int_{0}^{1}f_{\theta_{\star}}(x)\ln\left(\hbox{$\frac{1}{2}$}f_{\theta_{\star}}(x)\right)\mathrm{d}x$ $\leq\int_{0}^{1}f_{\theta_{\star}}(x)\left(\hbox{$\frac{1}{2}$}f_{\theta_{\star}}(x)-1\right)\mathrm{d}x$
$=\frac{1}{2}\int_{0}^{1}f_{\theta_{\star}}^{2}(x)-1$
$<0,$
for all $\theta_{\star}\geq 0$, or equivalently $\int_{0}^{1}f_{\theta_{\star}}(x)\ln f_{\theta_{\star}}(x)\,\mathrm{d}x<\ln 2$. Therefore, choosing $\delta>0$ small enough, we obtain
$\int_{0}^{1}K_{M}(x,\theta_{\star})f_{\theta_{\star}}(x)\,\mathrm{d}x\geq\ln(2-\delta)-c_{\theta_{\star}}>0$
for all $M>0$, a violation of assumption W3.
Appendix E Proof of Theorem 4
For all true data-generating parameters $\theta_{\star}\geq 0$, we prove that the MLE $\hat{\theta}_{n}$, if it exists, diverges to infinity a.s.-$F_{\theta_{\star}}^{\infty}$ in two steps as follows:
  1. 1.
    Fix a set of $n$ distinct numbers $\{x_{1},\dots,x_{n}\}\subset[0,1]$. We prove that, for any arbitrarily small $\delta>0$, there exists $\theta_{\delta}\geq 0$ such that
    $\frac{1+\cos(\theta_{\delta}x_{i})}{1+\frac{\sin\theta_{\delta}}{\theta_{\delta}}}\geq 2-\delta\quad\text{for all }i=1,\dots,n.$
    This immediately implies that, for all $n\in\mathbb{N}$, if $\hat{\theta}_{n}$ exists, then
    $\forall\delta>0,$ $\quad\prod_{i=1}^{n}f_{\hat{\theta}_{n}}(x_{i})\geq\prod_{i=1}^{n}f_{\theta_{\delta}}(x_{i})\geq(2-\delta)^{n}$
    $\implies$ $\quad\prod_{i=1}^{n}f_{\hat{\theta}_{n}}(x_{i})\geq 2^{n}$
  2. 2.
[…]
    $F_{\theta_{\star}}^{\infty}\left(x_{1:\infty}\in[0,1]^{\mathbb{N}}:\max_{\theta\in[0,M]}\prod_{i=1}^{n}f_{\theta}(x_{i})<2^{n}\quad\textnormal{ultimately}\right)=1$
    By step 1, this proves that the MLE is above $M$ for all large $n\in\mathbb{N}$ a.s.-$F_{\theta_{\star}}^{\infty}$, showing that (a) it is inconsistent at $\theta_{\star}$, and (b) it diverges to infinity a.s.-$F_{\theta_{\star}}^{\infty}$ (because $M$ is arbitrarily large). [^7] In step 2, any statement about the MLE $\hat{\theta}_{n}$ tacitly assumes its existence. Nevertheless, should it not exist for a certain finite sequence $y_{1:n}\in\mathbb{R}^{n}$ (i.e., should the supremum of the product likelihood not be achieved by any $\theta\geq 0$), any event $B\in\mathscr{B}(\mathbb{R}^{\mathbb{N}})$ of the form $B=\{x_{1:\infty}\in\mathbb{R}^{\mathbb{N}}:\hat{\theta}_{n}\in A\}$ (for some $A\in\mathscr{B}(\mathbb{R})$) should be understood to exclude all those infinite sequences $x_{1:\infty}\in\mathbb{R}^{\mathbb{N}}$ for which $x_{i}=y_{i}$ for all $i=1,...,n$.
E.1 Proof of step 1
Let $\delta>0$ be given, let $x_{1},\dots,x_{n}\in[0,1]$ be distinct numbers such that $x_{i}/\pi$ is irrational for all $i=1,\dots,n$, [^8] Notice that the set of all such configurations $(x_{1},\dots,x_{n})$ has $n$-dimensional Lebesgue measure 1, so restricting to such class of numbers is without loss of generality for our later probability statements. and assume that $\hat{\theta}_{n}$ exists. We wish to show that there exists $\theta>0$ such that
$\frac{1+\cos(\theta x_{i})}{1+\frac{\sin\theta}{\theta}}\geq 2-\delta\quad\text{for all }i=1,\dots,n.$ (7)
As for the denominator of Equation (7), note that
$\lim_{\theta\to\infty}\frac{\sin\theta}{\theta}=0.$
Thus, there exists $\theta_{0}\geq 0$ such that for all $\theta\geq\theta_{0}$ we have
$\left|\frac{\sin\theta}{\theta}\right|<\frac{\delta}{4}.$
In particular, for $\theta\geq\theta_{0}$,
$1+\frac{\sin\theta}{\theta}\leq 1+\frac{\delta}{4}.$
As for the numerator of Equation (7), we now want to find $\theta\geq\theta_{0}$ such that
$1+\cos(\theta x_{i})\geq 2-\frac{\delta}{2}\quad\text{for all }i=1,\dots,n,$
so it is enough to require
$\cos(\theta x_{i})\geq 1-\frac{\delta}{2}\quad\text{for all }i=1,\dots,n.$
By the continuity of the cosine function at 00, there exists $\varepsilon>0$ (depending on $\delta$) such that
$|\phi|<\varepsilon\quad\Longrightarrow\quad\cos\phi\geq 1-\frac{\delta}{2}.$
Thus, if we can find $\theta\geq\theta_{0}$ and integers $k_{1},\dots,k_{n}$ such that
$|\theta x_{i}-2\pi k_{i}|<\varepsilon\quad\text{for all }i=1,\dots,n,$
then we have
$\cos(\theta x_{i})=\cos\Bigl{(}\theta x_{i}-2\pi k_{i}\Bigr{)}\geq 1-\frac{\delta}{2},$
and consequently,
$1+\cos(\theta x_{i})\geq 2-\frac{\delta}{2}.$
To achieve this, we use Dirichlet’s simultaneous approximation theorem: [^9] See Corollary 1B on page 27 of Schmidt, (1980). for any irrational $x_{1}/2\pi,\dots,x_{n}/2\pi$ and for any $M>0$, there exist infinitely many natural numbers $\theta\geq M$ and integers $k_{1},\dots,k_{n}$ such that
$\left|\frac{x_{i}}{2\pi}-\frac{k_{i}}{\theta}\right|<\frac{1}{\theta^{1+1/n}}\quad\text{for all }i=1,\dots,n.$
Multiplying both sides of the inequality by $2\pi\theta$, we obtain
$|\theta x_{i}-2\pi k_{i}|<\frac{2\pi}{\theta^{1/n}}\quad\text{for all }i=1,\dots,n.$
Our goal is to have
$|\theta x_{i}-2\pi k_{i}|<\varepsilon.$
To ensure this, it suffices to have
$\frac{2\pi}{\theta^{1/n}}\leq\varepsilon\iff\theta\geq\left(\frac{2\pi}{\varepsilon}\right)^{n}.$
Hence, Dirichlet’s simultaneous approximation theorem guarantees the existence of a natural number $\theta$ and integers $k_{i}$ such that
$\theta\geq\max\left\{\theta_{0},\left(\frac{2\pi}{\varepsilon}\right)^{n}\right\}$
and
$|\theta x_{i}-2\pi k_{i}|<\varepsilon\quad\text{for all }i=1,\dots,n.$
In conclusion, for the integer $\theta$ obtained above we have $\theta\geq\theta_{0}$ and so
$1+\frac{\sin\theta}{\theta}\leq 1+\frac{\delta}{4}.$
Thus, for every $i=1,\dots,n$ we obtain
$\frac{1+\cos(\theta x_{i})}{1+\frac{\sin\theta}{\theta}}\geq\frac{2-\frac{\delta}{2}}{1+\frac{\delta}{4}}\geq 2-\delta.$
Remark 1.
As a byproduct of the above proof, we find that, for any $\delta>0$ and distinct points $(x_{1},\dots,x_{n})\in[0,1]^{n}$ in a set of $n$-dimensional full Lebesgue measure, there exist infinitely many $\theta>M$ (for any arbitrarily large $M>0$) at which the product likelihood takes a value greater than $(2-\delta)^{n}$. Because
$\max_{x\in[0,1]}f_{\theta}(x)=\frac{2\theta}{\theta+\sin\theta}\to 2\quad\textnormal{as }\theta\to\infty,$
this effectively means that, above any $M>0$, there is an infinite number of peaks of the likelihood whose height is arbitrarily close to the asymptotic maximum $2^{n}$.
E.2 Proof of step 2
In the second step of the proof, we show that the product likelihood, restricted to $\theta\in[0,M]$ for some $M<\infty$, cannot asymptotically attain the $2^{n}$ lower-bound derived in the previous step. To this end, we first establish that, for any fixed $\theta\geq 0$, the product of the likelihood values plus a small constant $\varepsilon>0$ cannot asymptotically reach this lower-bound. The presence of this positive $\varepsilon$ then allows us, together with the equicontinuity of the likelihood function, to extend the argument uniformly over $[0,M]$ via a standard covering argument. We therefore begin with the following lemma.
Lemma 2.
[…]
$\mathbb{E}\left[\prod_{i=1}^{n}(f_{\theta}(X_{i})+\varepsilon)\right]\leq(c+\varepsilon)^{n}$
for any $\theta,\varepsilon\geq 0$.
Proof.
By the iid assumption, we get
$\mathbb{E}\left[\prod_{i=1}^{n}(f_{\theta}(X_{i})+\varepsilon)\right]=\left(\int_{0}^{1}f_{\theta}(x)f_{\theta_{\star}}(x)\,\mathrm{d}x+\varepsilon\right)^{n}.$
Also
$\int_{0}^{1}f_{\theta}(x)$ $f_{\theta_{\star}}(x)\,\mathrm{d}x=\frac{\int_{0}^{1}(1+\cos(\theta x))(1+\cos(\theta_{\star}x))\mathrm{d}x}{\left(1+\frac{\sin\theta}{\theta}\right)\left(1+\frac{\sin\theta_{\star}}{\theta_{\star}}\right)}$
$=\frac{1+\frac{\sin\theta}{\theta}+\frac{\sin\theta_{\star}}{\theta_{\star}}+\frac{1}{2}\left[\int_{0}^{1}\cos((\theta-\theta_{\star})x)\mathrm{d}x+\int_{0}^{1}\cos((\theta+\theta_{\star})x)\mathrm{d}x\right]}{\left(1+\frac{\sin\theta}{\theta}\right)\left(1+\frac{\sin\theta_{\star}}{\theta_{\star}}\right)}$
$=\frac{1+\frac{\sin\theta}{\theta}+\frac{\sin\theta_{\star}}{\theta_{\star}}+\frac{1}{2}\left[\frac{\sin(\theta-\theta_{\star})}{\theta-\theta_{\star}}+\frac{\sin(\theta+\theta_{\star})}{\theta+\theta_{\star}}\right]}{\left(1+\frac{\sin\theta}{\theta}\right)\left(1+\frac{\sin\theta_{\star}}{\theta_{\star}}\right)}$
for all $\theta,\theta_{\star}\geq 0$. Therefore, letting $g(\theta):=\sin(\theta)/\theta$, we look for a constant $c<2$ satisfying the inequality
$c\geq\frac{1+2g(\theta)+\frac{1}{2}(1+g(2\theta))}{(1+g(\theta))^{2}}\equiv\frac{1+2g(\theta)+\frac{1}{2}(1+g(\theta)\cos(\theta))}{(1+g(\theta))^{2}}.$ (8)
for all $\theta\geq 0$, finding that $c=1.9$ works to this end. ∎
Now fix $\varepsilon\in(0,2-c)$ and $b\in(c+\varepsilon,2)$, where $c<2$ is the constant obtained in Lemma 2. Then, using Markov’s inequality, we obtain that
$F_{\theta_{\star}}^{\infty}\left(x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\,\prod_{i=1}^{n}(f_{\theta}(x_{i})+\varepsilon)\geq b^{n}\right)\leq\left(\frac{c+\varepsilon}{b}\right)^{n}\equiv e^{-dn},$ (9)
for all $\theta,\theta_{\star}\geq 0$ and $d\equiv\ln(b/(c+\varepsilon))>0$. Notice that $d$ itself, after fixing $\varepsilon$ and $b$, can be thought of as a universal constant independent of $\theta$ and $\theta_{\star}$. Now fix $M>0$, let $\{\theta^{1},\theta^{2},\dots\}$ be an enumeration of $\mathbb{Q}\cap[0,M]$, and define $\mathcal{T}_{n}=\{\theta^{1},\dots,\theta^{n}\}$. Therefore a union bound gives
$F_{\theta_{\star}}^{\infty}\left(x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\,\max_{\theta\in\mathcal{T}_{n}}\prod_{i=1}^{n}(f_{\theta}(x_{i})+\varepsilon)\geq b^{n}\right)\leq ne^{-dn}.$
Because this upper-bound is summable in $n\in\mathbb{N}$, the first Borel-Cantelli lemma yields $F_{\theta_{\star}}^{\infty}(\Omega_{\star})=1$, where
$\Omega_{\star}:=\left\{x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\,\max_{\theta\in\mathcal{T}_{n}}\prod_{i=1}^{n}(f_{\theta}(x_{i})+\varepsilon)<b^{n}\quad\textnormal{ultimately}\right\}.$
Now notice that, by density of $\mathbb{Q}\cap[0,M]$ in $[0,M]$, for all $\delta>0$ there exists $N_{\delta}\in\mathbb{N}$ such that, for all $n\geq N_{\delta}$, the maximum distance between consecutive points $\theta,\theta^{\prime}\in\mathcal{T}_{n}$ is less than $\delta$. In particular, fixing $n\geq N_{\delta}$, for all $\theta\in[0,M]$, there exists $\theta^{\prime}\in\mathcal{T}_{n}$ such that $|\theta-\theta^{\prime}|<\delta$. Moreover, by verifying that $|\partial f_{\theta}(x)/\partial\theta|\leq L$ for all $\theta\geq 0$, $x\in[0,1]$ and some $L<\infty$, the family $\{\theta\mapsto f_{\theta}(x):x\in[0,1]\}$ is easily seen to be equicontinuous. In particular, there exists $\delta>0$ such that $|\theta-\theta^{\prime}|<\delta$ implies $|f_{\theta}(x)-f_{\theta^{\prime}}(x)|<\varepsilon$ for all $x\in[0,1]$ and $\theta,\theta^{\prime}\in[0,M]$ (recall that $\varepsilon>0$ has been fixed beforehand).
Now fix $x_{1:\infty}\in\Omega_{\star}$. Therefore, there exists $N\in\mathbb{N}$ such that
$\max_{\theta\in\mathcal{T}_{n}}\prod_{i=1}^{n}(f_{\theta}(x_{i})+\varepsilon)<b^{n}$
for all $n\geq N$. Choosing $n\geq N\lor N_{\delta}$ and denoting $\theta_{n}\in\arg\max_{\theta\in[0,M]}\prod_{i=1}^{n}f_{\theta}(x_{i})$, [^10] Notice that, by an easy application of Weierstrass’s theorem, $\theta_{n}$ is well defined. we have that $|\theta_{n}-\theta|<\delta$ for some $\theta\in\mathcal{T}_{n}$, and so
$\prod_{i=1}^{n}f_{\theta_{n}}(x_{i})\leq\prod_{i=1}^{n}(f_{\theta}(x_{i})+\varepsilon)<b^{n}.$
Because this holds for any $x_{1:\infty}\in\Omega_{\star}$, we have shown that
$F_{\theta_{\star}}^{\infty}\left(x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\,\max_{\theta\in[0,M]}\prod_{i=1}^{n}f_{\theta}(x_{i})<b^{n}\quad\textnormal{ultimately}\right)=1.$
In particular, because $b<2$, by step 1 we have
$\left\{x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\,\max_{\theta\in[0,M]}\prod_{i=1}^{n}f_{\theta}(x_{i})<b^{n}\quad\textnormal{ultimately}\right\}\subseteq\left\{x_{1:\infty}\in[0,1]^{\mathbb{N}}\,:\hat{\theta}_{n}>M\quad\textnormal{ultimately}\right\},$
showing that the MLE $\hat{\theta}_{n}$ is larger than any $M>0$ for all large $n\in\mathbb{N}$, a.s.-$F_{\theta_{\star}}^{\infty}$.
Appendix F Proof of Proposition 3
Lemma 1 implies that the Hellinger metric entropy of the set $\Theta_{n}:=\{f_{\theta}:\theta\in[0,\bar{\theta}_{n}]\}$ can be bounded as follows. Because the Hellinger metric is upper-bounded by the Euclidean metric, one can cover $\Theta_{n}$ with $N\leq\bar{\theta}_{n}/\delta$ Euclidean, hence Hellinger, balls of radius $\delta$. Therefore the Hellinger $\delta$-covering number of $\Theta_{n}$ satisfies $N(\Theta_{n},d_{h},\delta)\leq\bar{\theta}_{n}/\delta$. In particular, this implies that the Hellinger metric entropy (the natural log of the covering number) satisfies
$\ln N(\Theta_{n},d_{h},\delta)\leq\ln\left(\frac{\bar{\theta}_{n}}{\delta}\right).$
Hence, we can use the conditions of Theorem 6.23 in Ghosal and van der Vaart, (2017) to ensure that the posterior mass of any Hellinger neighborhood of $f_{\theta_{\star}}$, for $\theta_{\star}=0$, converges to 1 a.s.-$F_{\theta_{\star}}^{\infty}$ as $n\to\infty$: for any $\delta>0$, we require that there exists $c>0$ such that, for all large $n\in\mathbb{N}$, the prior $\Pi$ satisfies $\Pi(\Theta_{n}^{c})\leq e^{-cn}$ as long as $\ln N(\Theta_{n},d_{h},\delta)\leq n\delta^{2}$, or equivalently $\bar{\theta}_{n}\leq\delta\exp\{n\delta^{2}\}$.
Therefore, any prior $\Pi$ on $[0,\infty)$ that, for all $\delta>0$, admits $c>0$ such that
$\Pi\left(\left[\delta\exp\{n\delta^{2}\},\,\infty\right)\right)\leq e^{-cn}\quad\textnormal{for all large }n\in\mathbb{N},$
will ensure Hellinger consistency at the uniform distribution. We now show that, for any $\varphi:(0,\infty)\to(0,\infty)$ such that $\lim_{t\to\infty}\varphi(t)/t=\infty$, a prior $\Pi$ with
$\Pi([\theta,\infty))\leq e^{-\varphi(\ln\theta)},\quad\textnormal{for all large }\theta>0$
satisfies the previous condition. Indeed, for all $\delta>0$, the properties of $\varphi$ imply that $\varphi(\ln\delta+n\delta^{2})\geq\ln\delta+n\delta^{2}\geq cn$ for all $n\in\mathbb{N}$ large and some small $c>0$. Therefore,
$\Pi\left(\left[\delta\exp\{n\delta^{2}\},\,\infty\right)\right)\leq\exp\left\{-\varphi(\ln\delta+n\delta^{2})\right\}\leq e^{-cn}\quad\textnormal{for all large }n\in\mathbb{N},$
as desired.
Appendix G Proof of Proposition 4
For any $\delta>0$, because the sequence $\theta_{k}:=(1+2k)\delta$, $k\in\mathbb{N}_{0}$, is a $\delta$-cover of $[0,\infty)$ (that is, for all $\theta\geq 0$, there exists $k\in\mathbb{N}_{0}$ such that $|\theta-\theta_{k}|\leq\delta$), the sequence $f_{\theta_{k}}$ is a $\delta$-cover, in the Hellinger sense, of $\{f_{\theta}:\theta\in[0,\infty)\}$ (thanks to Lemma 1). Therefore, defining
[…]
for all $k\in\mathbb{N}_{0}$, this shows that $A_{k}\subseteq A_{k}^{\star}:=\{f_{\theta}:d_{h}(f_{\theta},f_{\theta_{k}})\leq\delta\}$. Thus, we can use Theorem 4 in Walker, (2004) to design a prior $\Pi$ such that
[…]
and conclude that the posterior is Hellinger consistent at $\theta_{\star}=0$. In particular, we have assumed that $\Pi$ admits a density $\pi(\theta)$ on $[0,\infty)$ that, for all large $\theta>0$, is decreasing and satisfies
$\pi(\theta)\lesssim\frac{1}{\theta^{2}(\ln\theta)^{2+\beta}},$
for some $\beta>0$. Then, for $k_{0}\in\mathbb{N}_{0}$ large enough,
[…]
$=\frac{1}{\sqrt{2\delta}}\sum_{k=k_{0}}^{\infty}\frac{1}{k\ln(2k\delta)^{1+\beta/2}}<\infty$
for all $\delta>0$, showing that $\Pi$ yields strong consistency at $\theta_{\star}=0$.
Appendix H Proof of Theorem 5
For this proof, denote by $\hat{\theta}_{n}$ any MLE restricted to the sieve $[0,M_{n}]\cap A_{\varepsilon}^{c}$, where $M_{n}$ is a positive sequence determined throughout the proof, and $A_{\varepsilon}:=\{\theta\geq 0:d_{h}(f_{\theta},f_{0})<\varepsilon\}$ for some small $\varepsilon>0$. Then, for $\delta>0$ chosen small enough, for all $\theta\in A_{\varepsilon}^{c}$ we have
$F_{0}^{\infty}\left(\prod_{i=1}^{n}\left(f_{\theta}(X_{i})^{1/2}+\delta\right)\geq e^{-nb/2}\right)$ $\leq e^{nb/2}\left(\int_{0}^{1}\sqrt{f_{\theta}(x)f_{\theta_{\star}}(x)}\,\mathrm{d}x+\delta\right)^{n}$
$=e^{nb/2}\left(1-d_{h}^{2}(f_{\theta},f_{\theta_{\star}})/2+\delta\right)^{n}$
$\leq e^{nb/2}(1-\varepsilon^{2}/2+\delta)^{n}$
$\leq e^{nb/2}(1-\varepsilon^{2}/4)^{n}$
$\leq e^{nb/2}e^{-n\varepsilon^{2}/4},$
which, choosing $b>0$ small enough, is smaller than $e^{-nC_{\varepsilon}}$ for some $C_{\varepsilon}>0$. Let $M_{n}=e^{nc}$ for some $c<C_{\varepsilon}$ and let $\eta>0$ be such that $|\theta-\theta^{\prime}|<\eta\implies\max_{x\in[0,1]}|\sqrt{f_{\theta}(x)}-\sqrt{f_{\theta^{\prime}}(x)}|<\delta$ [^11] Recall that $\{\theta\mapsto f_{\theta}(x):x\in[0,1]\}$ is uniformly Lipschitz, therefore equicontinuous. Therefore, constructing an $\eta$-cover $\{\theta^{1},\theta^{2},\dots\}$ of $[0,M_{n}]\cap A_{\varepsilon}^{c}$ of cardinality at most $M_{n}/\eta$, a union bound gives
$F_{0}^{\infty}\left(\max_{\theta\in[0,M_{n}]\cap A_{\varepsilon}^{c}}\prod_{i=1}^{n}f_{\theta}^{1/2}(X_{i})\geq e^{-nb/2}\right)$ $\leq\frac{M_{n}}{\eta}F_{0}^{\infty}\left(\prod_{i=1}^{n}\left(f_{\theta^{1}}^{1/2}(X_{i})+\delta\right)\geq e^{-nb/2}\right)$
$\leq\frac{1}{\eta}e^{-(C_{\varepsilon}-c)n}.$
The above upper-bound is summable in $n\in\mathbb{N}$ and therefore we obtain that
$\prod_{i=1}^{n}\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\equiv\prod_{i=1}^{n}f_{\hat{\theta}_{n}}^{1/2}(X_{i})<e^{-nb/2}$ (10)
ultimately a.s.-$F_{0}^{\infty}$.
Now write
$\Pi(A_{\varepsilon}^{c}\mid X_{1:n})=\Pi(A_{\varepsilon}^{c}\cap[0,e^{cn}]\mid X_{1:n})+\Pi(A_{\varepsilon}^{c}\cap(e^{cn},\infty)\mid X_{1:n}).$ (11)
Using a line of reasoning similar to Walker and Hjort, (2001), rewrite the first addendum as follows:
$\Pi(A_{\varepsilon}^{c}\cap[0,e^{cn}]\mid X_{1:n})=\frac{\int_{A_{\varepsilon}^{c}\cap[0,e^{cn}]}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)}{\int_{\Theta}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)}.$ (12)
A standard result (see, e.g., Barron et al.,, 1999, Lemma 4) ensures that, as long as $\theta_{\star}\in\textnormal{KLS}(\Pi)$, the denominator satisfies
$\int_{\Theta}\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)>e^{-dn}$ (13)
ultimately a.s.-$F_{\theta_{\star}}^{\infty}$ for all $d>0$. For the numerator, we can write
$\int_{A_{\varepsilon}^{c}\cap[0,e^{cn}]}$ $\prod_{i=1}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\,\Pi(\mathrm{d}\theta)$
$\leq\prod_{i=1}^{n}\left(\frac{f_{\hat{\theta}_{n}}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\int_{A_{\varepsilon}^{c}\cap[0,e^{cn}]}\prod_{i=1}^{n}\left(\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\Pi(\mathrm{d}\theta).$ (14)
For the second factor in Equation (H), one obtains
$\mathbb{E}_{\theta_{\star}}\left[\int_{A_{\varepsilon}^{c}\cap[0,e^{cn}]}\prod_{i=1}^{n}\left(\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\Pi(\mathrm{d}\theta)\right]$ $\leq\int_{A_{\varepsilon}^{c}}\prod_{i=1}^{n}\mathbb{E}_{\theta_{\star}}\left[\left(\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\right]\Pi(\mathrm{d}\theta)$
$\leq\Pi(A_{\varepsilon}^{c})(1-\varepsilon^{2}/2)^{n}$
$\leq\Pi(A_{\varepsilon}^{c})e^{-n\varepsilon^{2}/2},$
so that, by Markov’s inequality and the first Borel–Cantelli lemma,
$\int_{A_{\varepsilon}^{c}\cap[0,e^{cn}]}\prod_{i=1}^{n}\left(\frac{f_{\theta}(X_{i})}{f_{\theta_{\star}}(X_{i})}\right)^{1/2}\Pi(\mathrm{d}\theta)<e^{-n\varepsilon^{2}/4}$ (15)
ultimately a.s.-$F_{\theta_{\star}}^{\infty}$. Thus, putting together Equations (10)-(15) and choosing $d>0$ small enough, we obtain
$\lim_{n\to\infty}\Pi(A_{\varepsilon}^{c}\cap[0,e^{cn}]\mid X_{1:n})=0$
a.s.-$F_{\theta_{\star}}^{\infty}$.
As for the second addendum in Equation (11), write
$\Pi(A_{\varepsilon}^{c}\cap(e^{cn},\infty)\mid X_{1:n})\leq\Pi((e^{cn},\infty)\mid X_{1:n})=\frac{\int_{e^{cn}}^{\infty}\prod_{i=1}^{n}f_{\theta}(X_{i})\pi(\theta)\mathrm{d}\theta}{\int_{0}^{\infty}\prod_{i=1}^{n}f_{\theta}(X_{i})\pi(\theta)\mathrm{d}\theta},$ (16)
and notice that, because the true density $f_{0}$ is equal to 1 on $[0,1]$, we can interpret $\prod_{i=1}^{n}f_{\theta}(X_{i})$ as the likelihood ratio. For the denominator, we once again invoke the KL support condition to satisfy Equation (13). As for the numerator, for all large $n\in\mathbb{N}$ we have
$\mathbb{E}\left[\int_{e^{cn}}^{\infty}\prod_{i=1}^{n}f_{\theta}(X_{i})\pi(\theta)\mathrm{d}\theta\right]$ $=\int_{e^{cn}}^{\infty}\mathbb{E}\left[\prod_{i=1}^{n}f_{\theta}(X_{i})\right]\pi(\theta)\mathrm{d}\theta$
$=\int_{e^{cn}}^{\infty}\pi(\theta)\mathrm{d}\theta$
$\lesssim\int_{e^{cn}}^{\infty}\theta^{-(1+\alpha)}\mathrm{d}\theta$
$=\frac{1}{\alpha}e^{-\alpha cn},$
[…]
$\lim_{n\to\infty}\Pi(A_{\varepsilon}^{c}\cap(e^{cn},\infty)\mid X_{1:n})=0$
a.s.-$F_{\theta_{\star}}^{\infty}$.